OpenStudy (anonymous):

The slope of a line is 2/3. What is the slope of a line that is perpendicular to this line?

OpenStudy (aravindg):

product of slopes will be -1 (2/3) * x=-1 solve for x

OpenStudy (anonymous):

So -1 is the anwser?

OpenStudy (aravindg):

nope solve for x

OpenStudy (aravindg):

x=-1/(2/3) =-1*3/2=-3/2

OpenStudy (anonymous):

Thanks just got that too

OpenStudy (aravindg):

yw :)